OpenStudy (yisselluna):

x + 5y = -2 2x + y = 5 The point of intersection of the lines has a y-coordinate of _____. -1 1 3

OpenStudy (kenrick143):

Do you know how to equate the lines?

OpenStudy (yisselluna):

no

OpenStudy (kenrick143):

Ok then, the lines touch one point so you could use x+5y=2 by equating.

OpenStudy (kenrick143):

I You can multiply it (x+5y=2) by 2 to eliminate x. And subtract it to 2x+y=5

OpenStudy (yisselluna):

i got that question rigjt already

OpenStudy (kenrick143):

Do what do you have now?

OpenStudy (kenrick143):

You got 9y=-9 right?

OpenStudy (kenrick143):

Got the answer?

OpenStudy (yisselluna):

yes